### v1

# Spectral properties of two-dimensional half-space semi-Dirac semi-metals

Abstract
The paper deals with the semi-Dirac operator in a half-space arising in the description of quasiparticles in quantum mechanics as well as in semi-metals materials and related structures. It completely shows the self-adjointness, computes the square and the spectrum of the operator. We also set up sufficient conditions for the existence of the point and discrete spectrum by including some appropriate potentials in the problem and study the spectral stability properties for perturbed operators.
Keywords Semi-Dirac semi-metals; Linear and quadratic dispersions; Quantum bound states.
1 Introduction
Semi-Dirac semi-metals are materials studied in the field of condensed matter physics with nanostructures can be featured as a combination of linear and quadratic dispersions and exhibiting tunable electronic, optical and transport properties. In spite of garnering substantial research attention, e.g. [3, 7, 11], the research of semi-Dirac semi-metals still remains a recent, fascinating and challenging topic in quantum materials.
Given a half-space domain  $\Omega=\mathbb{R}\times(0,\infty):=\mathbb{R}^{2}_{+}$ see Figure 1, on which the semi-Dirac semi-metals operator acts as
$T_{\mathbb{R}^{2}_{+}}:=\begin{pmatrix}-i\partial_{y}&-\partial^{2}_{x}+\delta\\ -\partial^{2}_{x}+\delta&i\partial_{y}\end{pmatrix}\qquad\mbox{in}\qquad L^{2}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})\,,$ (1)
here some physical constants are assumed to be equal to $1$ by virtue of re-scaling the system and the operator domain is defined as follows,
$\mathsf{D}(T_{\mathbb{R}^{2}_{+}})=\left\{u\in H^{1}(\mathbb{R}^{2}_{+};\mathbb{C}^{2}):\ \partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})\mbox{ and }u_{1}=u_{2}\ \mbox{on }y=0\right\}.$
Figure 1: The half-space +$\mathbb{R}^{2}_{+}$.
If the domain $\Omega=\mathbb{R}^{2}$ then the corresponding operator becomes
$T_{\mathbb{R}^{2}}:=\begin{pmatrix}-i\partial_{y}&-\partial^{2}_{x}+\delta\\ -\partial^{2}_{x}+\delta&i\partial_{y}\end{pmatrix}\qquad\mbox{in}\qquad L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})\,,$ (2)
$\mathsf{D}(T_{\mathbb{R}^{2}}):=\Big{\{}\psi=\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)\in H^{1}(\mathbb{R}^{2};\mathbb{C}^{2}):\ \partial_{x}^{2}\psi_{2},\partial_{x}^{2}\psi_{1}\in L^{2}(\mathbb{R}^{2})\Big{\}}$ (3)
As studied in [7], $T_{\mathbb{R}^{2}}$ is self-adjoint and its spectrum is purely absolutely continuous
$\sigma(T_{\mathbb{R}^{2}})=(-\infty,-\delta]\cup[\delta,\infty).$
Being motivated from [7], we study the self-adjointness, spectral properties of $T_{\mathbb{R}^{2}_{+}}$ and illustrate the existence of the discrete spectrum as well as the effect of some potentials on the spectrum of the semi-Dirac semi-metals problem.
The paper is organised as follows. In Section 2, we demonstrate the formula for the expectation value of the square of the operator. The self-adjointness of the semi-Dirac semi-metals operator in a half-space is established in Section 3. The results providing insights into the spectral properties of the operator, especially for the existence of the discrete spectrum and the spectral instability properties for perturbed operators are illustrated in Section 4.
2 The square of the semi-Dirac semi-metals operator in +$\mathbb{R}^{2}_{+}$
It is remarkable that $\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$ is a Banach space endowed with the norm
.$\|.\|_{T_{\mathbb{R}^{2}_{+};\mathbb{C}^{2}}}:=\sqrt{\|.\|^{2}_{H^{1}({\mathbb{R}^{2}_{+};\mathbb{C}^{2}})}+\|\partial_{x}^{2}.\|^{2}_{L^{2}({\mathbb{R}^{2}_{+};\mathbb{C}^{2}})}}\,.$
In addition, denote
.$\|.\|_{T_{\mathbb{R}^{2}_{+}}}:=\sqrt{\|.\|^{2}_{H^{1}({\mathbb{R}^{2}_{+};\mathbb{C}})}+\|\partial_{x}^{2}.\|^{2}_{L^{2}({\mathbb{R}^{2}_{+};\mathbb{C}})}}\,,$
$\mathtt{H}^{1}(\mathbb{R}^{2}_{+}):=\overline{C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C})}^{\|.\|_{T_{\mathbb{R}^{2}_{+}}}}\,,\quad\quad\quad\quad\,\,\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+}):=\overline{C_{0}^{\infty}(\mathbb{R}^{2}_{+};\mathbb{C})}^{\|.\|_{T_{\mathbb{R}^{2}_{+}}}}\,,$
$\mathtt{H}^{1}(\mathbb{R}^{2}_{+};\mathbb{C}^{2}):=\{u=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right),u_{1},u_{2}\in\mathtt{H}^{1}(\mathbb{R}^{2}_{+})\}\,,\quad\quad\mathtt{H}_{0}^{1}(\mathbb{R}^{2}_{+};\mathbb{C}^{2}):=\{u=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right),u_{1},u_{2}\in\mathtt{H}_{0}^{1}(\mathbb{R}^{2}_{+})\}\,.$
We have
$\mathtt{H}^{1}(\mathbb{R}^{2}_{+})=\{u\in H^{1}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}.$
The proof for this result can be analogously found in [1, 8]. To study the space $\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})$ thoroughly, we apply the approach and cut-off functions defined in [5].
Lemma 1.
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})=\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}$ (4)
Proof.
First of all, we easily check that
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})\subset\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}$
For the inverse inclusion, take a function $u\in H^{1}_{0}(\mathbb{R}^{2}_{+})$ such that $\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})$ then there exists a sequence $\{u_{n}\}_{n\geq 1}\subset C^{\infty}(\overline{\mathbb{R}^{2}_{+}})$: $u_{n}\xrightarrow[n\to\infty]{}u$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$. By virtue of the trace theorem, we have $u_{n}(x,0)\xrightarrow[n\to\infty]{}u(x,0)$ in $H^{\frac{1}{2}}(\mathbb{R})$.
For each $x\in\mathbb{R},y_{k}\in[0,\infty)$, we estimate
$|u_{n}(x,y_{k})|\leq|u_{n}(x,0)|+\int_{0}^{y_{k}}|\partial_{y}u_{n}(x,t)|\mathrm{d}t$ (5)
Hence,
$\int_{\mathbb{R}}|u_{n}(x,y_{k})|^{2}\mathrm{d}x\leq 2\Big{(}\int_{\mathbb{R}}|u_{n}(x,0)|^{2}\mathrm{d}x+y_{k}\int_{0}^{y_{k}}\int_{\mathbb{R}}|\partial_{y}u_{n}(x,t)|^{2}\mathrm{d}x\mathrm{d}t\Big{)}$ (6)
Letting $n\rightarrow\infty$, we have
$\int_{\mathbb{R}}|u(x,y_{k})|^{2}\mathrm{d}x\leq 2y_{k}\int_{0}^{y_{k}}\int_{\mathbb{R}}|\partial_{y}u(x,t)|^{2}\mathrm{d}x\mathrm{d}t$ (7)
for a.e. $y_{k}>0$.
Take $\varphi(y)\in C^{\infty}[0,\infty),0\leq\varphi\leq 1$
$\varphi(y):=\left\{\begin{array}[]{rcl}&1&\mbox{ if }y\in[0,1]\,,\\ &0&\mbox{ if }y\geq 2\,.\\ \end{array}\right.$
Putting
$\varphi_{n}(y):=\varphi(ny),\quad\quad\quad v_{n}:=u(x)(1-\varphi_{n})\,.$
Thus,
$\partial_{x}v_{n}=\partial_{x}u(1-\varphi_{n}),\quad\quad\quad\partial_{x}^{2}v_{n}=\partial_{x}^{2}u(1-\varphi_{n})$
and
$\partial_{y}v_{n}=\partial_{y}u(1-\varphi_{n})-u(\varphi_{n})^{\prime}\,.$
As a result,
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{x}v_{n}-\partial_{x}u|^{2}\leq\int_{\mathbb{R}^{2}_{+}}|\varphi_{n}|^{2}|\partial_{x}u|^{2}\mathrm{d}x\mathrm{d}y\xrightarrow[n\to\infty]{}0,$
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{x}^{2}v_{n}-\partial_{x}^{2}u|^{2}\leq\int_{\mathbb{R}^{2}_{+}}|\varphi_{n}|^{2}|\partial_{x}^{2}u|^{2}\mathrm{d}x\mathrm{d}y\xrightarrow[n\to\infty]{}0$
due to the dominated convergence theorem. Moreover, the support of $\varphi_{n}$, $\mathop{\mathrm{supp}}\nolimits\varphi_{n}\subset\mathbb{R}\times[0,\frac{2}{n}]$ and taking (7) into account, we have
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{y}v_{n}-\partial_{y}u|^{2}\mathrm{d}x\mathrm{d}y$ +$\leq\int_{\mathbb{R}^{2}_{+}}|\varphi_{n}|^{2}|\partial_{y}u|^{2}\mathrm{d}x\mathrm{d}y+\int_{0}^{\infty}\int_{\mathbb{R}}|u|^{2}|\varphi_{n}^{\prime}|^{2}\mathrm{d}x\mathrm{d}y\,,$ (8)
=+$=\int_{\mathbb{R}^{2}_{+}}|\varphi_{n}|^{2}|\partial_{y}u|^{2}\mathrm{d}x\mathrm{d}y+\int_{0}^{\frac{2}{n}}\int_{\mathbb{R}}n^{2}|u|^{2}|\varphi^{\prime}|^{2}\mathrm{d}x\mathrm{d}y\,,$
+$\leq\int_{\mathbb{R}^{2}_{+}}|\varphi_{n}|^{2}|\partial_{y}u|^{2}\mathrm{d}x\mathrm{d}y+2n^{2}\sup_{[0,\infty]}|\varphi^{\prime}|^{2}\Big{(}\int_{0}^{\frac{2}{n}}y\mathrm{d}y\Big{)}\Big{(}\int_{0}^{\frac{2}{n}}\int_{\mathbb{R}}|\partial_{y}u(x,y)|^{2}\mathrm{d}x\mathrm{d}y\Big{)}\,,$
$n\to\infty]{}0\,.$
Therefore, $v_{n}\xrightarrow[n\to\infty]{}u$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$. In addition, $v_{n}=0$ if $0\leq y\leq\frac{1}{n}$. Thus, we can modify the functions $u_{n}\in C^{\infty}_{0}(\mathbb{R}^{2}_{+})$ such that $u_{n}\xrightarrow[n\to\infty]{}u$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$. It deduces that $u\in\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})$. It concludes the proof.
[…]
Theorem 1.
For every $u=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right)\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$,
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$ (9)
Proof.
First of all, we will show that
$C^{\infty}(\mathbb{R}^{2})\cap\mathsf{D}(T_{\mathbb{R}^{2}_{+}})\mbox{ is a core of }T_{\mathbb{R}^{2}_{+}}.$ (10)
Recall that $\mathtt{H}^{1}(\mathbb{R}^{2}_{+}):=\overline{C^{\infty}(\overline{\mathbb{R}^{2}_{+}})}^{\|.\|_{T_{\mathbb{R}^{2}_{+}}}}$ then for each $u\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$ there exists a sequence $\{u_{1n}\}_{n\geq 1}\subset C^{\infty}(\overline{\mathbb{R}^{2}_{+}})$ such that $u_{1n}\xrightarrow[n\to\infty]{}u_{1}$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$.
Putting $v:=u_{2}-u_{1}$ then it follows that $v\in\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})$. Taking Lemma 1 into account, there exists a sequence $\{v_{n}\}_{n\geq 1}\in C^{\infty}_{0}(\mathbb{R}^{2}_{+})$ such that $v_{n}\xrightarrow[n\to\infty]{}v$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$. Hence, we deduce that $u_{1n}+v_{n}\xrightarrow[n\to\infty]{}u_{2}$ in $\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$. As a result, the smooth sequence $\{w_{n}:=\left(\begin{smallmatrix}u_{1n}\\ u_{1n}+v_{n}\end{smallmatrix}\right)\}_{n\geq 1}\xrightarrow[n\to\infty]{}u$ in $\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$. As a result, $T_{\mathbb{R}^{2}_{+}}w_{n}\xrightarrow[n\to\infty]{}T_{\mathbb{R}^{2}_{+}}u$ in $L^{2}(\mathbb{R}^{2}_{+})$. Therefore, (10) holds and thus we can compute the square of the operator acting on the smooth functions without loss of generality. We have
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}$ $=\|-i\partial_{y}u_{1}+(-\partial^{2}_{x}+\delta)u_{2}\|^{2}+\|(-\partial^{2}_{x}+\delta)u_{1}+i\partial_{y}u_{2}\|^{2}\,,$
$=\|-i\partial_{y}u_{1}\|^{2}+\|(-\partial^{2}_{x}+\delta)u_{2}\|^{2}+2\Re(-i\partial_{y}u_{1},(-\partial^{2}_{x}+\delta)u_{2})$
$\quad+\|(-\partial^{2}_{x}+\delta)u_{1}\|^{2}+\|i\partial_{y}u_{2}\|^{2}+2\Re((-\partial^{2}_{x}+\delta)u_{1},i\partial_{y}u_{2})\,,$
$=\|\nabla u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\Re(-\partial_{x}^{2}u_{1},u_{1})+2\delta\Re(-\partial_{x}^{2}u_{2},u_{2})$
$\quad+2\Re(-i\partial_{y}u_{1},(-\partial^{2}_{x}+\delta)u_{2})+2\Re((-\partial^{2}_{x}+\delta)u_{1},i\partial_{y}u_{2})\,,$
$=\|\nabla u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}$
$\quad+2\Re(-i\partial_{y}u_{1},(-\partial^{2}_{x}+\delta)u_{2})+2\Re((-\partial^{2}_{x}+\delta)u_{1},i\partial_{y}u_{2})\,.$
Regarding the boundary conditions and integration by parts, we estimate
$\Re(-i\partial_{y}u_{1},(-\partial^{2}_{x}+\delta)u_{2})$ =+$=-\Re\int_{\mathbb{R}^{2}_{+}}i\partial_{y}\overline{u_{1}}\partial^{2}_{x}u_{2}\ \mathrm{d}x\mathrm{d}y+\delta\Re\int_{\mathbb{R}^{2}_{+}}i\partial_{y}\overline{u_{1}}u_{2}\ \mathrm{d}x\mathrm{d}y\,,$
=+$=\Re\int_{\mathbb{R}^{2}_{+}}i\overline{u_{1}}\partial_{y}\partial_{x}^{2}u_{2}\mathrm{d}x\mathrm{d}y+\Re i\int_{\mathbb{R}}(\overline{u_{1}}\partial_{x}^{2}u_{2})|_{y=0}\mathrm{d}x$
$\quad-\delta\Re\int_{\mathbb{R}^{2}_{+}}i\overline{u_{1}}\partial_{y}u_{2}\ \mathrm{d}x\mathrm{d}y-\delta\Re\int_{\mathbb{R}}i(\overline{u_{1}}u_{2})|_{y=0}\mathrm{d}x\,,$
=+$=\Re\int_{\mathbb{R}^{2}_{+}}i\overline{u_{1}}\partial_{y}\partial_{x}^{2}u_{2}\mathrm{d}x\mathrm{d}y-\Re i\int_{\mathbb{R}}|\partial_{x}u_{2}|^{2}|_{y=0}\mathrm{d}x$
$\quad-\delta\Re\int_{\mathbb{R}^{2}_{+}}i\overline{u_{1}}\partial_{y}u_{2}\ \mathrm{d}x\mathrm{d}y+\delta\Re i\int_{\mathbb{R}}|u_{1}|_{y=0}\mathrm{d}x\,,$
=+$=\Re\int_{\mathbb{R}^{2}_{+}}i\partial_{x}^{2}\overline{u_{1}}\partial_{y}u_{2}\mathrm{d}x\mathrm{d}y-\delta\Re\int_{\mathbb{R}^{2}_{+}}i\overline{u_{1}}\partial_{y}u_{2}\ \mathrm{d}x\mathrm{d}y\,,$
$=-\Re((-\partial^{2}_{x}+\delta)u_{1},i\partial_{y}u_{2})\,.$
Combining all previous computations, we get the desired formula. This completes the proof.
Remark 1.
For other domains, e.g. rectangular domains, the density (10) may not hold due to the fact that the validity of Lemma 1 is not true for general sets.
3 The self-adjointness of the semi-Dirac semi-metals operator
First of all, we study on the closeness of $T_{\mathbb{R}^{2}_{+}}$. The result can be directly gained by using the definition of the operator closeness.
Proposition 1.
$T_{\mathbb{R}^{2}_{+}}$ is a closed operator.
Proof.
Take a sequence $\{u_{n}\}_{n\geq 1}\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}}):u_{n}\xrightarrow[n\to\infty]{}u$ and $T_{\mathbb{R}^{2}_{+}}u_{n}\xrightarrow[n\to\infty]{}v$. Then we have $u_{n}\xrightarrow[n\to\infty]{}u\in H^{1}(\mathbb{R}^{2}_{+})$ and there exists $w\in L^{2}(\mathbb{R}^{2}_{+})$ such that $\partial_{x}^{2}u_{n}\xrightarrow[n\to\infty]{}w$ in $L^{2}(\mathbb{R}^{2}_{+})$. In addition, for all $\varphi\in C^{\infty}_{0}(\mathbb{R}^{2}_{+})$
$\left\langle u,\partial_{x}^{2}\varphi\right\rangle=\lim_{n\rightarrow\infty}\left\langle u_{n},\partial_{x}^{2}\varphi\right\rangle=\lim_{n\rightarrow\infty}\left\langle\partial_{x}^{2}u_{n},\varphi\right\rangle=\left\langle w,\varphi\right\rangle\,,$ (11)
[…]
Recall that $T_{\mathbb{R}^{2}}$ is self-adjoint. To point out the self-adjointness of $T_{\mathbb{R}^{2}_{+}}$, we employ the extension map as the approach described in [2]. We will establish the trace continuous maps
Putting
$T:=\begin{pmatrix}-i\partial_{y}&-\partial^{2}_{x}+\delta\\ -\partial^{2}_{x}+\delta&i\partial_{y}\end{pmatrix}\,,$ (12)
acts as a differential operator with distribution sense.
Lemma 2.
Let :=$\mathcal{M}:=\{u\in L^{2}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})|Tu\in L^{2}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})\}$ endowed with the norm
$\|u\|_{\mathcal{M}}^{2}:=\|u\|_{L^{2}(\mathbb{R}^{2}_{+})}^{2}+\|Tu\|_{L^{2}(\mathbb{R}^{2}_{+})}^{2}$
then $\mathcal{M}$ is a Hilbert space and $C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$ is dense in $\mathcal{M}$.
Proof.
The proof is similar as that for the Dirac operators, see [4]. Indeed, it is obvious that $\mathcal{M}$ is a Hilbert space encoded with a scalar product
$(u,v)_{\mathcal{M}}=(u,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,Tv)_{L^{2}(\mathbb{R}^{2}_{+})}$
To prove $C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$ is dense in $\mathcal{M}$, we take $f\in\mathcal{M}$ such that $(f,g)_{\mathcal{M}}=0$ for all $g\in C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$. Putting $h:=Tg$ then $Th=-f$ in the sense of distributions. Denote $\tilde{f},\tilde{h}$ are the trivial extensions of $f,h$ to $\mathbb{R}^{2}$. For all $\phi\in C^{\infty}_{0}(\mathbb{R}^{2};\mathbb{C}^{2})$ we have
$(\tilde{h},T\phi)_{L^{2}(\mathbb{R}^{2})}=(h,T\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-f,\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-\tilde{f},\phi)_{L^{2}(\mathbb{R}^{2})}.$
It yields that $T\tilde{h}=-\tilde{f}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$. Therefore, $\tilde{h}\in H^{1}(\mathbb{R}^{2};\mathbb{C}^{2}),\partial_{x}^{2}\tilde{h}\in L^{2}(\mathbb{R}^{2})$ by means of the Fourier transform. As a result, $h\in H^{1}(\mathbb{R}^{2}_{+};\mathbb{C}^{2}),\partial^{2}_{x}h\in L^{2}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})$ and $h=0$ on $y=0$ due to the trace theorem. Equivalently, $h\in\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})$ then there exists a sequence $h_{n}\in C^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{C}^{2}):h_{n}\xrightarrow[n\to\infty]{}h$ in $\mathtt{H}_{0}^{1}(\mathbb{R}^{2}_{+};\mathbb{C}^{2})$.
For all $u\in\mathcal{M}$, we have
$(f,u)_{\mathcal{M}}$ $=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(h,Tu)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$ (13)
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(h_{n},Tu)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(Th_{n},u)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(Th,u)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
$=0\,.$
Hence, $f=0$ and thus, $C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$ is dense in $\mathcal{M}$. The proof is completed.
In order to establish a continuous trace map for $\mathcal{M}$, we recall the definition of Fourier transform of a function $f(t)\in L^{2}(\mathbb{R}^{N})$
$\hat{f}(\xi)=\int_{\mathbb{R}^{N}}e^{-ix\xi}f(t)\mathrm{d}t\,,$
and its inverse transform
$f(t)=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{ix\xi}\hat{f}(\xi)\mathrm{d}\xi\,.$
We have the Plancherel identity, see [6]
$\int_{\mathbb{R}^{N}}|\hat{f}(\xi)|^{2}\mathrm{d}\xi=(2\pi)^{N}\int_{\mathbb{R}^{N}}|f(x)|^{2}\mathrm{d}x\,.$
Putting
$\mathtt{H}^{\frac{1}{2}}:=\{u(x)\in L^{2}(\mathbb{R};\mathbb{C}):\int_{\mathbb{R}}|\hat{u}(\xi_{1})|^{2}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\mathrm{d}\xi_{1}:=\|u\|^{2}_{\mathtt{H}^{\frac{1}{2}}}<\infty\}.$
Then $\mathtt{H}^{\frac{1}{2}}$ is also a Hilbert space and its dual space is
$\mathtt{H}^{-\frac{1}{2}}:=\{u(x)\in L^{2}(\mathbb{R};\mathbb{C}):\int_{\mathbb{R}}|\hat{u}(\xi_{1})|^{2}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{-\frac{1}{2}}\mathrm{d}\xi_{1}:=\|u\|^{2}_{\mathtt{H}^{-\frac{1}{2}}}<\infty\}.$
Denote
$\mathtt{H}^{\frac{1}{2}}_{\mathbb{C}^{2}}:=\{u(x)=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right),u_{1},u_{2}\in\mathtt{H}^{\frac{1}{2}}\}\,,$
$\mathtt{H}^{-\frac{1}{2}}_{\mathbb{C}^{2}}:=\{u(x)=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right),u_{1},u_{2}\in\mathtt{H}^{-\frac{1}{2}}\}\,,$
we have the following theorem
Theorem 2.
For every $u\in\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$, we have $u(x,0)\in\mathtt{H}^{\frac{1}{2}}$ and $\|u(x,0)\|_{\mathtt{H}^{\frac{1}{2}}}\leq C\|u\|_{\mathtt{H}^{1}(\mathbb{R}^{2}_{+})}$ with $C$ is a constant. Reversely, for each $g\in\mathtt{H}^{\frac{1}{2}}$ there exists a function $u\in\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$ and a constant $M>0$ such that $g=u(x,0)$ and $\|u\|_{\mathtt{H}^{1}(\mathbb{R}^{2}_{+})}\leq M\|u(x,0)\|_{\mathtt{H}^{\frac{1}{2}}}$.
Proof.
For the first statement, we use similar arguments for Sobolev spaces as described in [6] to prove the theorem, we can suppose $u\in C^{\infty}(\overline{\mathbb{R}^{2}_{+}})$ without loss of generality due to the density of smooth functions up to the boundary in $\mathtt{H^{1}}(\mathbb{R}^{2}_{+})$. First of all, we put $g(x):=u(x,0)$, extend $u$ to $\mathbb{R}^{2}$
$\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}u(x,y)&\mbox{ if }y\geq 0\,,\\ u(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$ (14)
then $\tilde{u}\in H^{1}(\mathbb{R}^{2})$ and $\partial_{x}^{2}\tilde{u}\in L^{2}(\mathbb{R}^{2})$. Denote $\hat{u}$ is the Fourier transform of $\tilde{u}$, we deduce that
$\hat{g}(\xi_{1})=\frac{1}{2\pi}\int_{\mathbb{R}}\hat{u}(\xi)\mathrm{d}\xi_{2}$
and
$\int_{\mathbb{R}}|\hat{g}(\xi_{1})|^{2}$ $(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\mathrm{d}\xi_{1}=\frac{1}{4\pi^{2}}\int_{\mathbb{R}}\mid\int_{\mathbb{R}}\hat{u}(\xi)\mathrm{d}\xi_{2}\mid^{2}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\mathrm{d}\xi_{1}\,,$
$\leq\frac{1}{4\pi^{2}}\int_{\mathbb{R}}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\int_{\mathbb{R}}|\hat{u}(\xi)|^{2}(1+|\xi|^{2}+|\xi_{1}|^{4})\mathrm{d}\xi_{2}\int_{\mathbb{R}}(1+|\xi|^{2}+|\xi_{1}|^{4})^{-1}\mathrm{d}\xi_{2}\mathrm{d}\xi_{1}\,.$
Moreover,
$\int_{\mathbb{R}}(1+|\xi|^{2}+|\xi_{1}|^{4})^{-1}\mathrm{d}\xi_{2}=\frac{\pi}{(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}}.$
Thus,
$\int_{\mathbb{R}}|\hat{g}(\xi_{1})|^{2}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\mathrm{d}\xi_{1}\leq\frac{1}{4\pi}\int_{\mathbb{R}^{2}}|\hat{u}(\xi)|^{2}(1+|\xi|^{2}+|\xi_{1}|^{4})\mathrm{d}\xi.$
Equivalently, we obtain
$\|u(x,0)\|_{\mathtt{H}^{\frac{1}{2}}}\leq C\|u\|_{\mathtt{H}^{1}(\mathbb{R}^{2}_{+})}.$
Reversely, for each $g\in\mathtt{H}^{\frac{1}{2}}$ we construct a function
$\hat{u}(\xi):=2\,\frac{\hat{g_{1}}(\xi_{1})}{\xi_{2}^{2}+(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})}\,,$
where
$\hat{g_{1}}(\xi_{1})=\hat{g}(\xi_{1})(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\,.$
We have
$\int_{\mathbb{R}^{2}}|\hat{u}(\xi)|^{2}(1+|\xi|^{2}+|\xi_{1}|^{4})\mathrm{d}\xi$ $=4\int_{\mathbb{R}^{2}}\frac{|\hat{g_{1}}(\xi_{1})|^{2}}{\xi_{2}^{2}+1+|\xi_{1}|^{2}+|\xi_{1}|^{4}}\mathrm{d}\xi_{1}\mathrm{d}\xi_{2}\,,$
$\leq 4\int_{\mathbb{R}}|\hat{g_{1}}(\xi_{1})|^{2}\frac{\pi}{(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}}\mathrm{d}\xi_{1}\,,$
$\leq 4\pi\int_{\mathbb{R}}|\hat{g}(\xi_{1})|^{2}(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})^{\frac{1}{2}}\mathrm{d}\xi_{1}\,,$
$\leq 4\pi\|g\|^{2}_{\mathtt{H}^{\frac{1}{2}}}\,.$
Therefore, $u(x,y)\in\mathtt{H}^{1}(\mathbb{R}^{2}_{+})$ for $y\geq 0$ and and $2\sqrt{\pi}\|g(x)\|_{\mathtt{H}^{\frac{1}{2}}}\geq 2\pi\|u\|_{\mathtt{H}^{1}(\mathbb{R}^{2}_{+})}$.
Moreover, we take $u_{n}\in C^{\infty}(\overline{\mathbb{R}^{2}};\mathbb{C}^{2}):u_{n}\xrightarrow[n\to\infty]{}u$ in $\mathsf{D}(T_{\mathbb{R}^{2}})$ then $\gamma u_{n}(x,0)\xrightarrow[n\to\infty]{}\gamma u(x,0)$ in $\mathtt{H}^{\frac{1}{2}}(\mathbb{R})$ and $\hat{u}_{n}(\xi)\xrightarrow[n\to\infty]{}\hat{u}(\xi)$ in $L^{2}(\mathbb{R}^{2})$. Using the inverse Fourier transform, we estimate
$u_{n}(x,y)=\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{ix\xi_{1}+iy\xi_{2}}\hat{u}_{n}(\xi)\mathrm{d}\xi.$
Thus,
$u_{n}(x,0)$ $=\frac{1}{4\pi^{2}}\int_{\mathbb{R}^{2}}e^{ix\xi_{1}}\hat{u}_{n}(\xi)\mathrm{d}\xi\,,$ (15)
$n\to\infty]{}\frac{1}{4\pi^{2}}\int_{\mathbb{R}^{2}}e^{ix\xi_{1}}\hat{u}(\xi)\mathrm{d}\xi=\frac{1}{2\pi^{2}}\int_{\mathbb{R}^{2}}e^{ix\xi_{1}}\frac{\hat{g_{1}}(\xi_{1})}{\xi_{2}^{2}+(1+|\xi_{1}|^{2}+|\xi_{1}|^{4})}\mathrm{d}\xi\,,$
$=\frac{1}{2\pi}\int_{\mathbb{R}}e^{ix\xi_{1}}\hat{g}(\xi_{1})\mathrm{d}\xi_{1}=g(x)\,.$
Therefore, $\gamma u(x,0)=g(x)\in\mathtt{H}^{\frac{1}{2}}(\mathbb{R})$. It concludes the proof. ∎
Theorem 3.
There exists a continuous trace map $\gamma:\mathcal{M}\rightarrow\mathtt{H}^{-\frac{1}{2}}$.
Proof.
For each $u\in\mathcal{M}$, there exists a sequence $\{u_{n}\}_{n\geq 1}\in C^{\infty}(\overline{\mathbb{R}^{2}_{+}})$ such that $u_{n}\xrightarrow[n\to\infty]{}u$ in $\mathcal{M}$ due to Lemma 2. We will show that the traces $\gamma u_{n}$ converges in $\mathtt{H}_{\mathbb{C}^{2}}^{-\frac{1}{2}}$.
Take $g=\left(\begin{smallmatrix}g_{1}\\ g_{2}\end{smallmatrix}\right)\in\mathtt{H}_{\mathbb{C}^{2}}^{\frac{1}{2}}$, there exists a function $v=\left(\begin{smallmatrix}v_{1}\\ v_{2}\end{smallmatrix}\right)\in\mathtt{H}_{\mathbb{C}^{2}}^{1}$ such that $\|v\|_{\mathtt{H}_{\mathbb{C}^{2}}^{1}}\leq 4\|g\|_{\mathtt{H}_{\mathbb{C}^{2}}^{\frac{1}{2}}},\gamma v=g$ by virtue of Theorem 2. We estimate
$(u_{n},Tv)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1n}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2n}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$ (16)
$=(Tu_{n},v)+\int_{\mathbb{R}}i\overline{u_{1n}}v_{1}|_{y=0}-i\overline{u_{2n}}v_{2}|_{y=0}\ \mathrm{d}x\,.$
Therefore,
$(u_{n}-u_{m},Tv)$ $=(Tu_{n}-Tu_{m},v)+\int_{\mathbb{R}}i\overline{(u_{1n}-u_{1m})}v_{1}|_{y=0}-i\overline{(u_{2n}-u_{2m})}v_{2}|_{y=0}\ \mathrm{d}x$ (17)
for all $m,n\geq 1$.
Now we choose $g_{2}=v_{2}=0$. It deduces that
$|(u_{1n}(x,0)-u_{1m}(x,0),v)|$ $\leq\|u_{m}-u_{n}\|\|Tv\|+\|Tu_{n}-Tu_{m}\|\|v\|\,,$
$\quad\leq C(\|u_{n}-u_{m}\|+\|Tu_{n}-Tu_{m}\|)\|v\|_{\mathtt{H}_{\mathbb{C}^{2}}^{1}}\,,$
$\quad\leq M(\|u_{n}-u_{m}\|+\|Tu_{n}-Tu_{m}\|)\|v\|_{\mathtt{H}_{\mathbb{C}^{2}}^{\frac{1}{2}}}\,,$
with $M,C$ are constants being independent of $v$ and $\|Tv\|$ is defined by virtue of Theorem 1. Hence, $\lim_{n\rightarrow\infty}(u_{1n}-u_{1m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. Analogously, if we choose $g_{1}=v_{1}=0$, we also obtain $\lim_{n\rightarrow\infty}(u_{2n}-u_{2m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. As a result, there exists a trace map
$\gamma:\quad\quad$ $\mathcal{M}\longrightarrow\mathtt{H}_{\mathbb{C}^{2}}^{-\frac{1}{2}}$
$u\longmapsto\gamma u=\lim_{n\rightarrow\infty}u_{n}(x,0)\,.$
The limit is independent of the chosen convergent sequences. It means that if $\{u_{n}\}_{n\geq 1},\{u^{\prime}_{n}\}_{n\geq 1}$ simultaneously converge to $u$ in $\mathcal{M}$ then $\gamma u$ is still uniquely identified.
Theorem 4.
$T_{\mathbb{R}^{2}_{+}}$ is a self-adjoint operator.
Proof.
For all $u,v\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$, we have
$(u,T_{\mathbb{R}^{2}_{+}}v)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$ (18)
$=\int_{0}^{\infty}\int_{\mathbb{R}}i\partial_{y}\overline{u_{1}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{u_{1}}v_{2}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}i\overline{u_{1}}v_{1}|_{y=0}\ \mathrm{d}x$
$\quad+\int_{0}^{\infty}\int_{\mathbb{R}}-i\partial_{y}\overline{u_{2}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{u_{2}}v_{1}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}-i\overline{u_{2}}v_{2}|_{y=0}\ \mathrm{d}x\,,$
$=(T_{\mathbb{R}^{2}_{+}}u,v)\,.$
It implies that $T_{\mathbb{R}^{2}_{+}}$ is symmetric due to the fact that $u_{1}=u_{2},v_{1}=v_{2}$ on $y=0$.
Now we take $u\in\mathsf{D}(T^{*}_{\mathbb{R}^{2}_{+}})$, recall the definition for the domain of the adjoint operator
$\mathsf{D}(T^{*}_{\mathbb{R}^{2}_{+}})$ :=$:=\left\{u\in L^{2}(\mathbb{R}^{2}_{+})\mbox{ such that}\ \forall v\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})\mbox{ there exists }\phi_{u}\in L^{2}(\mathbb{R}^{2}_{+}):(u,T_{\mathbb{R}^{2}_{+}}v)=(\phi_{u},v)\right\}\,.$ (19)
For each $w=\left(\begin{smallmatrix}w_{1}\\ w_{2}\end{smallmatrix}\right)$, we denote $w^{*}:=\left(\begin{smallmatrix}w_{2}\\ w_{1}\end{smallmatrix}\right)$. Taking $v\in C^{\infty}_{0}(\mathbb{R}^{2}_{+})$ we have
$(u,T_{\mathbb{R}^{2}_{+}}v)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$ (20)
$=(Tu,v)\leq\|\phi_{u}\|_{L^{2}(\mathbb{R}^{2}_{+})}\|v\|_{L^{2}(\mathbb{R}^{2}_{+})}\,.$
It deduces that $Tu\in L^{2}(\mathbb{R}^{2}_{+})$ and thus, $Tu=\phi_{u}$.
For all $v\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$, we have
$(u,T_{\mathbb{R}^{2}_{+}}v)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$ (21)
$=\int_{0}^{\infty}\int_{\mathbb{R}}i\partial_{y}\overline{u_{1}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{u_{1}}v_{2}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}i\overline{u_{1}}v_{1}|_{y=0}\ \mathrm{d}x$
$\quad+\int_{0}^{\infty}\int_{\mathbb{R}}-i\partial_{y}\overline{u_{2}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{u_{2}}v_{1}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}-i\overline{u_{2}}v_{2}|_{y=0}\ \mathrm{d}x\,,$
$=(Tu,v)+\int_{\mathbb{R}}i\overline{(u_{1}-u_{2})}v_{1}|_{y=0}\ \mathrm{d}x\,.$
It yields that $u_{1}=u_{2}$ in $\mathtt{H}^{-\frac{1}{2}}$.
Now we extend the functions $u$ to $\mathbb{R}^{2}$
$\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}u(x,y)&\mbox{ if }y\geq 0\,,\\ u(x,-y)^{*}&\mbox{ if }y<0\,,\\ \end{array}\right.$ (22)
For all $v\in C^{\infty}(\mathbb{R}^{2};\mathbb{C}^{2})$, we put $w(x,y)=v(x,-y)$ for all $(x,y)\notin\mathbb{R}^{2}_{+}$ then we deduce that $w\in C^{\infty}(\mathbb{R}^{2}\setminus\mathbb{R}^{2}_{+})$. Taking $u_{1}=u_{2}$ in $\mathtt{H}^{-\frac{1}{2}}$ into account, we have
$(\tilde{u},Tv)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{-\infty}^{0}\int_{\mathbb{R}}\overline{\tilde{u}_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{\tilde{u}_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}i\partial_{y}\overline{\tilde{u}_{1}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{1}}v_{2}\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{\mathbb{R}}-i\overline{\tilde{u}_{1}}v_{1}|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}-i\partial_{y}\overline{\tilde{u}_{2}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{2}}v_{1}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}i\overline{\tilde{u}_{2}}v_{2}|_{y=0}\ \mathrm{d}x\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+((Tu)^{*},w)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}-i(\overline{u_{2}}v_{1}-\overline{u}_{1}v_{2})|_{y=0}\ \mathrm{d}x\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,w^{*})_{L^{2}(\mathbb{R}^{2}_{+})}\leq\|Tu\|_{L^{2}(\mathbb{R}^{2}_{+})}\|v\|_{L^{2}(\mathbb{R}^{2})}\,.$
It implies that $\tilde{u}\in\mathsf{D}(T^{*}_{\mathbb{R}^{2}})$. Moreover, $\mathsf{D}(T^{*}_{\mathbb{R}^{2}})=\mathsf{D}(T_{\mathbb{R}^{2}})$ due to the self-adjointess of $T_{\mathbb{R}^{2}}$. Hence, we deduce that $u\in H^{1}(\mathbb{R}^{2}_{+})$ and $\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})$. Besides, we also have $u^{*}(x,0)=u(x,0)$ due to the trace theorem and thus, it deduces that $u_{1}=u_{2}$ on $y=0$. This concludes the proof of the theorem.
4 The spectral properties in semi-Dirac semi-metals
By virtue of the self-adjointness, the closest-to-zero squared spectrum of the operator $T_{\mathbb{R}^{2}_{+}}$ can be computed by
$\lambda_{1}(\mathbb{R}^{2}_{+})^{2}=\inf_{\stackrel{{\scriptstyle[}}{{u}}\not=0]{}{u\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})}}\frac{\|T_{\mathbb{R}^{2}_{+}}u\|^{2}}{\,\|u\|^{2}}$
where
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$
Hence,
$\lambda_{1}(\mathbb{R}^{2}_{+})^{2}-\delta^{2}=\inf_{\stackrel{{\scriptstyle[}}{{u}}\not=0]{}{u\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})}}\frac{|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}}{\,\|u\|^{2}}\geq 0$
Thus, the spectrum of $T_{\mathbb{R}^{2}_{+}}$: $\sigma_{T_{\mathbb{R}^{2}_{+}}}\subset(-\infty,-\delta]\cup[\delta,\infty)$. Moreover, this spectrum is purely essential, we have the following lemma
Lemma 3.
The point spectrum of $T_{\mathbb{R}^{2}_{+}}$ is empty.
Proof.
Suppose that the point spectrum of $T_{\mathbb{R}^{2}_{+}}$ is not empty then there exists $\lambda\in\mathbb{R},u=\left(\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right)\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$ being the eigenvalue and eigenfunction of $T_{\mathbb{R}^{2}_{+}}$ respectively, we have
{$-i\partial_{y}u_{1}+(-\partial_{x}^{2}+\delta)u_{2}&=\lambda u_{1}\quad&&\mbox{in}\quad\mathbb{R}^{2}_{+}\,,\\ (-\partial_{x}^{2}+\delta)u_{1}+i\partial_{y}u_{2}&=\lambda u_{2}\quad&&\mbox{in}\quad\mathbb{R}^{2}_{+}\,,\\ u_{1}&=u_{2}\quad&&\mbox{on }y=0.\end{aligned}\right.$ (23)
Extending $u$ to $\mathbb{R}^{2}$, we put
$\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}u(x,y)&\mbox{ if }y\geq 0\,,\\ u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$ (24)
It yields
$\partial_{x}\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}\partial_{x}u(x,y)&\mbox{ if }y\geq 0\,,\\ \partial_{x}u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.\quad\quad\quad\partial_{x}^{2}\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}\partial_{x}^{2}u(x,y)&\mbox{ if }y\geq 0\,,\\ \partial_{x}^{2}u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$
and
$\partial_{y}\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}\partial_{y}u(x,y)&\mbox{ if }y\geq 0\,,\\ -\partial_{y}u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$
It deduces that $\tilde{u}\in H^{1}(\mathbb{R}^{2}),\partial^{2}_{x}\tilde{u}\in L^{2}(\mathbb{R}^{2})$ and $\tilde{u}$ is an eigenfunction corresponding to an eigenvalue $\lambda$ of $T_{\mathbb{R}^{2}}$. As a result, the point spectrum of $T_{\mathbb{R}^{2}}$ is not empty. It is a contrary to the fact that the spectrum of $T_{\mathbb{R}^{2}}$ is purely absolutely continuous and the point spectrum of $T_{\mathbb{R}^{2}}$ is empty by means of the Fourier transform. Consequently, the point spectrum of $T_{\mathbb{R}^{2}_{+}}$ is empty. The proof is completed.
Regarding Lemma 3 and the self-adjointness of $T_{\mathbb{R}^{2}_{+}}$, it is apparent that the spectrum of $T_{\mathbb{R}^{2}_{+}}$ is purely essential : $\sigma(T_{\mathbb{R}^{2}_{+}})=\sigma_{\textup{ess}}(T_{\mathbb{R}^{2}_{+}})$. Now we compute the spectrum of $T_{\mathbb{R}^{2}_{+}}$ as follows,
Theorem 5.
$\sigma_{T_{\mathbb{R}^{2}_{+}}}=(-\infty,-\delta]\cup[\delta,\infty).$
Proof.
Recall that $\sigma_{T_{\mathbb{R}^{2}_{+}}}\subset(-\infty,-\delta]\cup[\delta,\infty)$ due to the min-max principle. Therefore, we only have to verify the opposite.
Take $k\in\mathbb{R},k\geq 0$ and choose a function $\varphi(x,y)\in C^{\infty}_{0}(\mathbb{R}^{2}_{+})$ such that $\|\varphi\|=\frac{1}{\sqrt{2}}$. Putting
$\varphi_{n}(x,y):=\frac{1}{n}\varphi(\frac{x}{n},\frac{y}{n})$
then we also have $\|\varphi_{n}\|=\frac{1}{\sqrt{2}}$. Moreover,
$\partial_{x}\varphi_{n}(x,y)=\frac{1}{n^{2}}\partial_{x}\varphi(\frac{x}{n},\frac{y}{n}),\quad\quad\quad\partial_{y}\varphi_{n}(x,y)=\frac{1}{n^{2}}\partial_{y}\varphi(\frac{x}{n},\frac{y}{n})\,,$
$\partial_{x}^{2}\varphi_{n}(x,y)=\frac{1}{n^{3}}\partial_{x}^{2}\varphi(\frac{x}{n},\frac{y}{n})\,.$
Therefore,
$\|\partial_{x}\varphi_{n}(x,y)\|=\frac{1}{n}\|\partial_{x}\varphi\|,\quad\|\partial_{y}\varphi_{n}(x,y)\|=\frac{1}{n}\|\partial_{y}\varphi\|,\quad\|\partial_{x}^{2}\varphi_{n}(x,y)\|=\frac{1}{n^{2}}\|\partial_{x}^{2}\varphi\|\,.$
Putting
$u_{n}(x,y):=\varphi_{n}(x,y)e^{ikx},\quad\quad\psi_{n}:=\left(\begin{smallmatrix}u_{n}\\ u_{n}\end{smallmatrix}\right)$
then $\psi_{n}\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}}),\|\psi_{n}\|=1$ and
$\partial_{x}u_{n}=\partial_{x}\varphi_{n}e^{ikx}+ik\varphi_{n}e^{ikx},\quad\quad\partial_{x}^{2}u_{n}=\partial_{x}^{2}\varphi_{n}e^{ikx}+2ik\,\partial_{x}\varphi_{n}e^{ikx}-k^{2}u_{n}\,,$
$\partial_{y}u_{n}=\partial_{y}\varphi_{n}e^{ikx}\,.$
For every $\mu\in[\delta,\infty)$, we choose $k^{2}=\mu-\delta$ and estimate
$\|T_{\mathbb{R}^{2}_{+}}\psi_{n}-\mu\psi_{n}\|^{2}$ $=\|-i\partial_{y}u_{n}-\mu u_{n}+(-\partial_{x}^{2}+\delta)u_{n}\|^{2}+\|i\partial_{y}u_{n}-\mu u_{n}+(-\partial_{x}^{2}+\delta)u_{n}\|^{2}\,,$ (25)
$=\|-i\partial_{y}u_{n}-\partial_{x}^{2}\varphi_{n}e^{ikx}-2ik\,\partial_{x}\varphi_{n}e^{ikx}\|^{2}+\|i\partial_{y}u_{n}-\partial_{x}^{2}\varphi_{n}e^{ikx}-2ik\,\partial_{x}\varphi_{n}e^{ikx}\|^{2}\,,$
$\leq 2(\|\partial_{y}\varphi_{n}\|^{2}+\|\partial_{x}^{2}\varphi_{n}\|^{2}+\|2k\,\partial_{x}\varphi_{n}\|^{2})\xrightarrow[n\to\infty]{}0\,.$
It follows that $\mu\in\sigma(T_{\mathbb{R}^{2}_{+}})$ due to the Weyl theorem.
On the other hand, for every $\mu\in(-\infty,-\delta]$, we choose $k^{2}=-\delta-\mu$ and $\phi_{n}=\left(\begin{smallmatrix}u_{n}\\ -u_{n}\end{smallmatrix}\right)\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$. Employing the analogous estimates, we also obtain that
$\|T_{\mathbb{R}^{2}_{+}}\phi_{n}-\mu\phi_{n}\|\xrightarrow[n\to\infty]{}0\,.$
Equivalently, $\mu\in\sigma(T_{\mathbb{R}^{2}_{+}})$. Hence, it concludes the proof.
Summarily, we have investigated that $\sigma(T_{\mathbb{R}^{2}_{+}})=\sigma_{\textup{ess}}(T_{\mathbb{R}^{2}_{+}})=(-\infty,-\delta]\cup[\delta,\infty)$. In order to establish the point spectrum of the semi-Dirac semi-metals, we need to add some potentials.
Let us take $V\in L^{\infty}(\mathbb{R}^{2}_{+};\mathbb{R})$, we consider the following operator
$H:=T_{\mathbb{R}^{2}_{+}}+V\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,\mbox{ with }\mathsf{D}(H):=\mathsf{D}(T_{\mathbb{R}^{2}_{+}})\,.$ (26)
By the boundedness of $V$ and the self-adjointness of $T_{\mathbb{R}^{2}_{+}}$ then it is apparent that $H$ is also self-adjoint. Now we illustrate the effect of the sign of $V$ on the properties of the spectrum of $H$.
Proposition 2.
If $L^{\infty}(\mathbb{R}^{2}_{+};\mathbb{R})\ni V=V(x)\geq 0$ a.e. in +$\mathbb{R}^{2}_{+}$ then $\sigma(H)\subset(-\infty,-\delta]\cup[\delta,\infty)$.
Proof.
By the hypothesis, for any $u\in\mathsf{D}(H)$, we have
$\|Hu\|^{2}$ $=\|\Big{(}T_{\mathbb{R}^{2}_{+}}+V\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\Big{)}u\|^{2}$
$=\|-i\partial_{y}u_{1}+(-\partial^{2}_{x}+\delta+V)u_{2}\|^{2}+\|(-\partial^{2}_{x}+\delta+V)u_{1}+i\partial_{y}u_{2}\|^{2}\,,$
$=\|-i\partial_{y}u_{1}\|^{2}+\|(-\partial^{2}_{x}+\delta+V)u_{2}\|^{2}+2\Re(-i\partial_{y}u_{1},(-\partial^{2}_{x}+\delta+V)u_{2})$
$\quad+\|(-\partial^{2}_{x}+\delta+V)u_{1}\|^{2}+\|i\partial_{y}u_{2}\|^{2}+2\Re((-\partial^{2}_{x}+\delta+V)u_{1},i\partial_{y}u_{2})\,,$
$=\|\partial_{y}u\|^{2}+\|(-\partial^{2}_{x}+\delta+V)u\|^{2}\geq\delta^{2}\|u\|^{2}\,.$
It follows that $\sigma(H)\subset(-\infty,-\delta]\cup[\delta,\infty)$ due to the min-max principle. This completes the proof.
Theorem 6.
There exists $V=V(x,y)\in L^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{R})$ such that the discrete spectrum of $H$ is not empty.
Proof.
First of all, it is striking that $\sigma_{\textup{ess}}(H)=\sigma_{\textup{ess}}(T_{\mathbb{R}^{2}_{+}})=(-\infty,-\delta]\cup[\delta,\infty)$ due to the boundedness and vanishing at infinity of $V$ as discussed in [7, 9, 10].
We study on the Dirichlet Laplacian defined on a segment $[a,b]$ with $0<a<b$. Denote $\lambda_{1},u_{1}(x)$ and $\lambda_{1},u_{2}(y)$ are the ground-state eigenvalues and eigenfunctions of the Dirichlet Laplacian with respect to variables $x,y$ determined on $[a,b]$ such that $\|u_{1}\|_{L^{2}([a,b])}=\|u_{2}(y)\|_{L^{2}([a,b])}=1$. As a result, $\lambda_{1}=\frac{\pi^{2}}{(b-a)^{2}}$. Thoroughly, $u_{1}(x),u_{2}(y)$ satisfy the following system with respect to $z=x$ and $z=y$
$-\partial_{z}^{2}\psi(z)&=\lambda_{1}\psi(z)&&\mbox{in}&&[a,b]\,,\\ \psi(a)&=\psi(b)=0\,.\end{aligned}\right.$ (27)
Putting $\psi(x,y):=u_{1}(x)u_{2}(y),v:=\left(\begin{smallmatrix}\psi\\ \psi\end{smallmatrix}\right)$. In order to reduce notations, we can identity $v$ with the trivial extension of $v$ to $\mathbb{R}^{2}$ and thus, we have $v\in\mathsf{D}(H)$. It is obvious that $\|\psi\|=1,\|v\|=\sqrt{2}$.
Now we choose $V\in C^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{R})$ such that $V$ is a constant in $[a,b]\times[a,b]$, we compute
$\|Hv\|^{2}$ $=2\|\partial_{y}\psi\|^{2}+2\|(-\partial_{x}^{2}+\delta+V)\psi\|^{2}\,,$ (28)
$=2\|u_{1}(x)\|_{L^{2}(a,b)}^{2}\|\partial_{y}u_{2}\|_{L^{2}(a,b)}^{2}+2\|u_{2}(y)\|^{2}_{L^{2}(a,b)}\|(\lambda_{1}+\delta+V)u_{1}\|_{L^{2}(a,b)}^{2}\,,$
$=2\lambda_{1}+2\|(\lambda_{1}+\delta+V)u_{1}\|_{L^{2}(a,b)}^{2}\,.$
It yields that
$\|Hv\|^{2}-\delta^{2}\|v\|^{2}$ $=2\lambda_{1}+2\lambda_{1}^{2}+4\delta\lambda_{1}+4\lambda_{1}V+4\delta V+2V^{2}\ \,,$ (29)
$=2V^{2}+4(\lambda_{1}+\delta)V+2\lambda_{1}+4\delta\lambda_{1}+2\lambda_{1}^{2}\,.$
The reduced discriminant of this quadratic trinomial equals
$4(\lambda_{1}+\delta)^{2}-4\lambda_{1}-8\lambda_{1}\delta-4\lambda_{1}^{2}=4\delta^{2}-4\lambda_{1}.$
If $\lambda_{1}=\frac{\pi^{2}}{(b-a)^{2}}<\delta^{2}$ then the corresponding quadratic equation has two solutions $V_{1},V_{2}:V_{1}<V_{2}$. Therefore, for any $V\in(V_{1},V_{2})$ in $[a,b]^{2}$ we have $\|Hv\|^{2}-\delta^{2}\|v\|^{2}<0$. It deduces the existence of the discrete spectrum of $H^{2}$ as well as $H$. It concludes the proof.
Now we establish perturbed operators defined in [7] to analyse the spectral stability in small perturbations, we denote
$W:=\begin{pmatrix}w_{11}&w_{12}\\ w_{21}&w_{22}\end{pmatrix}\,,\quad\quad H_{\epsilon}:=T_{\mathbb{R}^{2}_{+}}+\epsilon W,,$
where $w_{12},w_{21}:\mathbb{R}^{2}_{+}\rightarrow\mathbb{C}$ in $L^{\infty}(\mathbb{R}^{2}_{+})$, $w_{ii}\in L^{\infty}(\mathbb{R}^{2}_{+};\mathbb{R})$ for $i=1,2$, $w_{ij}$ vanishes at infinity for every $i,j=1,2$ and $\mathsf{D}(H_{\epsilon}):=\mathsf{D}(T_{\mathbb{R}^{2}_{+}})$.
Recall that if $w_{12}=\overline{w_{21}}$ then $H_{\epsilon}$ is self-adjoint [7] and $\sigma_{\textup{ess}}(H_{\epsilon})=\sigma_{\textup{ess}}(T_{\mathbb{R}^{2}_{+}})=(-\infty,-\delta]\cup[\delta,\infty)$ due to the boundedness and vanishing at infinity of $W$. This result can be seen as the stability of the essential spectrum for the perturbed operators.
Considering a sequence of approximate functions of $1$ in $L^{2}(\mathbb{R}^{2}_{+})$ as analogously defined in [7], we choose $g(t)\in C^{\infty}([0,1];\mathbb{R}):0\leq g\leq 1\,\forall t\in[0,1]$,
$g(t)=\left\{\begin{array}[]{rcl}0&\mbox{ if }t\in[0,\frac{1}{9}]\,,\\ 1&\mbox{ if }t\in[\frac{1}{2},1]\,,\\ \end{array}\right.$
and $g_{n}(x,y)\in C^{\infty}(\mathbb{R}^{2}_{+}),n\geq 1$
$g_{n}(x,y)=\left\{\begin{array}[]{rcl}&1&\mbox{ if }\sqrt{x^{2}+y^{2}}\leq n\,,\\ &g(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})&\mbox{ if }n<\sqrt{x^{2}+y^{2}}<n^{2}\,,\\ &0&\mbox{ if }\sqrt{x^{2}+y^{2}}\geq n^{2}\,,\\ \end{array}\right.$ (30)
Lemma 4.
[…]
Proof.
The pointwise convergence is obvious due to the definition of $g_{n}$.
Compute partial derivatives of $g_{n}$, we have
$\partial_{x}g_{n}(x,y)=\left\{\begin{array}[]{rcl}&0&\mbox{ if }\sqrt{x^{2}+y^{2}}\in[0,n]\cup[n^{2},\infty)\,,\\ &-\frac{1}{\ln n}\,\frac{x}{x^{2}+y^{2}}\,g^{\prime}(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})&\mbox{ if }\sqrt{x^{2}+y^{2}}\in(n,n^{2})\,,\\ \end{array}\right.$
$\partial_{y}g_{n}(x,y)=\left\{\begin{array}[]{rcl}&0&\mbox{ if }\sqrt{x^{2}+y^{2}}\in[0,n]\cup[n^{2},\infty)\,,\\ &-\frac{1}{\ln n}\,\frac{y}{x^{2}+y^{2}}\,g^{\prime}(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})&\mbox{ if }\sqrt{x^{2}+y^{2}}\in(n,n^{2})\,,\\ \end{array}\right.$
$\partial_{x}^{2}g_{n}(x,y)=\left\{\begin{array}[]{rcl}&0&\mbox{ if }\sqrt{x^{2}+y^{2}}\in[0,n]\cup[n^{2},\infty)\,,\\ &f(x,y)&\mbox{ if }\sqrt{x^{2}+y^{2}}\in(n,n^{2})\,,\\ \end{array}\right.$
where
$f(x,y)=\frac{1}{\ln^{2}n}\,\frac{x^{2}}{(x^{2}+y^{2})^{2}}\,g^{\prime\prime}(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})+\frac{1}{\ln n}\,\frac{x^{2}-y^{2}}{(x^{2}+y^{2})^{2}}\,g^{\prime}(\frac{\ln\frac{n^{2}}{\sqrt{x^{2}+y^{2}}}}{\ln n})\,.$
Passing to polar coordinates, we estimate
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{x}g_{n}|^{2}\mathrm{d}x\mathrm{d}y$ $=\int_{0}^{\pi}\int_{n}^{n^{2}}\frac{\cos^{2}\varphi}{r^{2}\ln^{2}n}|g^{\prime}(\frac{\ln n^{2}-\ln r}{\ln n})|^{2}r\mathrm{d}r\mathrm{d}\varphi\,,$
$=\frac{\pi}{2}\frac{1}{\ln n}\int_{0}^{1}|g^{\prime}(t)|^{2}\mathrm{d}t\xrightarrow[n\to\infty]{}0\,,$
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{y}g_{n}|^{2}\mathrm{d}x\mathrm{d}y$ $=\int_{0}^{\pi}\int_{n}^{n^{2}}\frac{\sin^{2}\varphi}{r^{2}\ln^{2}n}|g^{\prime}(\frac{\ln n^{2}-\ln r}{\ln n})|^{2}r\mathrm{d}r\mathrm{d}\varphi\,,$
$=\frac{\pi}{2}\frac{1}{\ln n}\int_{0}^{1}|g^{\prime}(t)|^{2}\mathrm{d}t\xrightarrow[n\to\infty]{}0\,.$
+$\int_{\mathbb{R}^{2}_{+}}|\partial_{x}^{2}g_{n}|^{2}\mathrm{d}x\mathrm{d}y$ $\leq 2\int_{0}^{\pi}\int_{n}^{n^{2}}\frac{\cos^{4}\varphi}{r^{4}\ln^{4}n}|g^{\prime\prime}(\frac{\ln n^{2}-\ln r}{\ln n})|^{2}r\mathrm{d}r\mathrm{d}\varphi\,$
$\quad+2\int_{0}^{\pi}\int_{n}^{n^{2}}\frac{\cos^{2}2\varphi}{r^{4}\ln^{2}n}|g^{\prime}(\frac{\ln n^{2}-\ln r}{\ln n})|^{2}r\mathrm{d}r\mathrm{d}\varphi\,,$
$=\frac{3\pi}{4}\frac{1}{n^{2}\ln^{3}n}\int_{0}^{1}|g^{\prime\prime}(t)|^{2}\mathrm{d}t+\frac{\pi}{n^{2}\ln n}\int_{0}^{1}|g^{\prime}(t)|^{2}\mathrm{d}t\xrightarrow[n\to\infty]{}0\,.$
Hence, the proof is completed.
Putting
$A_{\epsilon}:=\int_{\mathbb{R}^{2}_{+}}(\epsilon^{2}w_{11}^{2}+\epsilon^{2}w_{12}^{2}+4\delta\epsilon\Re w_{12}+\epsilon^{2}w_{21}^{2}+\epsilon^{2}w_{22})\,\mathrm{d}x\mathrm{d}y\,,$
we consider the existence of the discrete spectrum not lying in the essential spectrum of $H_{\epsilon}$ depending on the sign of $A_{\epsilon}$.
Theorem 7.
Let $w_{ij}\in L^{1}(\mathbb{R}^{2}_{+})\cap L^{\infty}(\mathbb{R}^{2}_{+})$ for each $i,j=1,2$ and vanishing at infinity. If $A_{\epsilon}<0$ then the discrete spectrum in $-(\delta,\delta)$ of $H_{\epsilon}$ is non-empty.
Proof.
Denote
$\psi_{n}:=\left(\begin{smallmatrix}g_{n}(x,y)\\ g_{n}(x,y)\end{smallmatrix}\right)\,,$
here $g_{n}$ is previously defined in (30). Thus, it deduces that $\psi_{n}\in\mathsf{D}(H_{\epsilon})$, we have
$\|H\psi_{n}\|^{2}-\delta^{2}\|\psi_{n}\|^{2}$ $=\|(-i\partial_{y}+\epsilon w_{11})g_{n}+(-\partial^{2}_{x}+\delta+\epsilon w_{12})g_{n}\|^{2}$ (31)
$\quad+\|(-\partial^{2}_{x}+\delta+\epsilon w_{21})g_{n}+(i\partial_{y}+\epsilon w_{22})g_{n}\|^{2}\,,$
$n\to\infty]{}\int_{\mathbb{R}^{2}_{+}}(\epsilon^{2}w_{11}^{2}+\epsilon^{2}w_{12}^{2}+4\delta\epsilon\Re w_{12}+\epsilon^{2}w_{21}^{2}+\epsilon^{2}w_{22})\,\mathrm{d}x\mathrm{d}y\,,$
due to the dominated convergence theorem and Lemma 4.
If $A_{\epsilon}<0$ then there exist $n*\in\mathbb{N},B_{\epsilon}<0$ such that for all $*n\geq n*$ : $\|H\psi_{n}\|^{2}-\delta^{2}\|\psi_{n}\|^{2}<B_{\epsilon}<0$. As a result, the discrete spectrum of $H_{\epsilon}$ is non-empty or there exists at least one eigenvalue in $(-\delta,\delta)$ of $H_{\epsilon}$. It concludes the proof.
Theorem 7 demonstrates the spectral instability for a sequence of operators $H_{\epsilon}$ when there always exists an electromagnetic potential $W$ such that $\sigma(H_{\epsilon})\neq\sigma(H)=\sigma(H_{\epsilon=0})$.
Another comment referring to the case $w_{11}=w_{22}=0,w_{12}=w_{21}\geq 0$ is if there exists the discrete spectrum of $H_{\epsilon}$ then that has to be embedded eigenvalues lying in the essential spectrum of $H_{\epsilon}$. This result is derived by using the analogous arguments given in [7].
Corollary 1.
If +$\int_{\mathbb{R}^{2}_{+}}\Re w_{12}\ \mathrm{d}x\mathrm{d}y<0$ then there exists $\epsilon(\delta)>0$ such that for all $\epsilon\in(0,\epsilon(\delta))$ the discrete spectrum of $H_{\epsilon}$ is non-empty.
Proof.
The proof is directly deduced from Theorem 7. Here we can choose
$\epsilon(\delta):=\frac{-4\delta\int_{\mathbb{R}^{2}_{+}}\Re w_{12}}{\|w_{11}\|^{2}+\|w_{22}\|^{2}+\|w_{12}\|^{2}+\|w_{21}\|^{2}}\,.$
It is apparent that
$\lim_{\delta\rightarrow 0}\epsilon(\delta)=0,\quad\quad\quad\quad\quad\lim_{\delta\rightarrow\infty}\epsilon(\delta)=\infty\,.$
The proof is completed.
It is noteworthy that the sufficient conditions posed on $\Re w_{12}$ in Corollary 1 and in Theorem 6 are not identical. The former relies on the integral of $\Re w_{12}$ in +$\mathbb{R}^{2}_{+}$, while the latter requires rigorous values of $V=\Re w_{12}$ in a certain range $[a,b]^{2}\in\mathbb{R}^{2}_{+}$.
Acknowledgment
The author was partially supported by the EXPRO grant No. 20-17749X of the Czech Science Foundation.
Declarations
Conflict of interest There is no conflict of interest to declare.